OpenStudy (anonymous):

Will Medal And Fan Best Answer!! I need explanation AND answer! Divide. Write your answer in simplest form. 9/5 divided by 5/2

OpenStudy (anonymous):

oh this is so easy one sec...

OpenStudy (anonymous):

Ok so when you divide fraction, you need to always apply KCF

Nnesha (nnesha):

\[\huge\rm \frac{ \frac{ a }{ b } }{ \frac{ c }{ d } } =\frac{ a }{ b } \times \frac{ d }{ c}\] multiply top fraction with the reciprocal of the bottom fraction

OpenStudy (anonymous):

the answer is 18/25 and ill explain now why

Nnesha (nnesha):

first step) explain 2nd) let them answer it. your teacher will never allow you to check answer first frm the book

OpenStudy (anonymous):

So you first have 9/5 divided by 5/2. You have to do kcf (keep, change ,flip) whenever you have to divide . So keep 9/5 the same. CHANGE the sign from division to multiplication and FLIP the fraction 5/2 to 2/5. Now your equation is 9/5 times 2/5. 9 times 2 is 18 and 5 times 5 is 25... Take a look at the drawing.